a shoe store sold 340 pairs of shoes in one day. The number of pairs of athletic shoes sold was 4 more than twice the number of pairs of dress shoes sold. How many pairs of athletic shoes were sold?
Answer by nyc_function(2741)


x+y=340
x+4=2y
x=2y-4
substitute for x
(2y-4)+y=340
3y=336
y=112
x=116
116 pair of athletic shoes and 112 pair of dress shoes
I hope this helps.